OpenStudy (anonymous):

Integral(sqroot(1+sin(theta))

OpenStudy (anonymous):

i'm missing a rule just not sure which one

OpenStudy (anonymous):

click on "show steps" on wolfram...theres some crazy U subs that only a highly trained mind would see

OpenStudy (anonymous):

what you got is not correct the solution manual says -cos(x)/sqroot(sin(x)-1))

OpenStudy (anonymous):

\[\int \sqrt{1+\sin(\theta)}d\theta\]

OpenStudy (anonymous):

sure it is correct, just a different form.

OpenStudy (anonymous):

yes that is the integral only it has other parts let me put the whole thing here\[\int\limits_{0}^{2\pi}\sqrt{2}(\sqrt{1+\sin(\theta)}d \theta\]

OpenStudy (anonymous):

its -(2/3(cos theta))*(1 +sin theta)^(3/2) thats the right one

OpenStudy (anonymous):

How to confuse everyone in one easy lesson...